OpenStudy (anonymous):

Help please ??? Ill medal Bo and Erica are yoga instructors. Between the two of them, they teach 35 yoga classes each week. If Erica teaches 10 fewer than twice as many as Bo, how many classes does each instructor teach per week? A 12 Bo; 23 Erica B 19 Bo; 16 Erica C 17 Bo; 18 Erica D 15 Bo; 20 Erica

OpenStudy (anonymous):

between the 2 of them, they teach 35 classes B + E = 35 Eric teaches 10 fewer then twice as many as Bo E = 2B - 10 you can now solve by substitution....do you know how to do that ??

OpenStudy (anonymous):

no , show me

OpenStudy (anonymous):

u will use the first eq to solve the second one

OpenStudy (anonymous):

B + E = 35 B = 35 - E now sub 35 - E in for B back into the other equation E = 2B - 10 E = 2(35 - E) - 10 E = 70 - 2E - 10 3E = 60 E = 20 now sub 20 in for E and solve for B B + E = 35 B + 20 = 35 B = 35 - 20 B = 15 so Bo (B) teaches 15 and Erica (E) teaches 20 questions ?

OpenStudy (anonymous):

wow very good explanation